OpenStudy (anonymous):

solve 3x^-2+x^-1-10=0

OpenStudy (anonymous):

i got x=1/2 and x=-3/5

OpenStudy (anonymous):

jut not sure if that came out correctly

OpenStudy (lgbasallote):

simplify by turning the exponents to positive by putting them in the denom 3/x^2 + 1/x - 10 = 0 howd you get those may i ask?

OpenStudy (anonymous):

3x^-2+x^-1-10=0